OpenStudy (anonymous):

Solve these two equations : 2sin^2 x - 5sin x - 3 = 0 (cos x - (square root of 2)/2)(sec x - 1) = 0

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

@aelong

OpenStudy (anonymous):

is someone online?

OpenStudy (wolf1728):

Hello QuenBee - wow Open Study is acting so nutty it took me a few minutes just to click in here.

OpenStudy (anonymous):

Let's sue OS XDD

OpenStudy (anonymous):

haha ok could you guys help me out with those two questions please?

OpenStudy (wolf1728):

I think I have solved the first equation x = 330°

OpenStudy (anonymous):

lol! can you explain?

OpenStudy (anonymous):

sorry last time I gave an answer like that to my teacher (x= so and so) and nothing else she laughed at my face and told me to do the assignment again

OpenStudy (wolf1728):

Well, sorry for giving an answer. I guess you want to see the work too huh?

OpenStudy (anonymous):

so far I have sin x = -1/2 and sin x = 3?

OpenStudy (anonymous):

no, no you misunderstood I meant thanks but an explanation would be great

OpenStudy (anonymous):

so could you show the steps on how u got there?

OpenStudy (wolf1728):

2sin^2 x - 5sin x - 3 = 0 Actually, I programmed the equation into an electronic spread sheet and found it by trial and error.

OpenStudy (anonymous):

oh

OpenStudy (wolf1728):

I imagine the equation has to be manipulated algebraically, etc?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I have (2sin x + 1)(sin x -3) = 0

OpenStudy (wolf1728):

2 sin² (x) -6sin(x) +sin(x) -3 =0 THat seems to be factored correctly.

OpenStudy (anonymous):

2 sin^2(x) - 6sin(x) - 3 is the one that's correct

OpenStudy (anonymous):

?

OpenStudy (wolf1728):

No, I just multiplied the factoring that you got. Basically, I'm saying you are right.

OpenStudy (wolf1728):

Multiplying (2sin x + 1)(sin x -3) equals 2 sin² (x) -6sin(x) +sin(x) -3 =0

OpenStudy (anonymous):

so for x I get sin^-1(-1/2) and 3 is that correct?

OpenStudy (anonymous):

sin^-1 (-1/2) would be pi/6?

OpenStudy (anonymous):

hey u there?

OpenStudy (anonymous):

hello??